You lift heavy book from the floor of the room and keep it in the book - shelf having height 2m. In this process you take 5 seconds. The work done by you will depend upon
 

Grade:11

1 Answers

Khimraj
5 years ago
work done = energy*t = mgh*t = m*10*2*5 = 100m
where m is mass of book.